OpenStudy (djjo):

Please help! Will give medal. Algebra 1

OpenStudy (djjo):

Can someone help me with C?

OpenStudy (djjo):

I don't know what to write/answer

OpenStudy (anonymous):

so set x as the years, set width as y, for western beach y=100-2x, and for dunes beach y=20+5x

OpenStudy (anonymous):

Solve them as simultaneous equations

OpenStudy (djjo):

Thank you @14mdaz!!!